# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

### v2

@@ II-C2 Problem
TABLE VI: Properties and methods of problem base class.
@@ D Number of decision variable dimensions for all tasks
getOptimum() Get optimum solutions for multi-objective optimization
@@ 11  Prob = Prob@Problem(varargin);
All problems within MToP inherit from the problem base class Problem.m. The properties and methods of Problem.m are detailed in Table VI. As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach [19]. For task $k$ with upper and lower bounds $L_{k}$ and $U_{k}$, respectively, and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is linearly mapped as follows:
$\bm{x}^{\prime}=\frac{\bm{x}-L_{k}}{U_{k}-L_{k}}.$ (2)
The dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$. During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision. It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
@@
A simple problem implementation example is illustrated in Listing LABEL:list:prob. The maximum number of function evaluations, denoted as maxFE, for the problem can be specified within the class constructor. The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task. Within this function, the number of tasks T is first defined, followed by a detailed setup for each task. Subsequently, parameters such as decision variable dimensions D, objective dimensions M, fitness function Fnc, lower bound Lb, and upper bound Ub are set individually for each task. For problems involving multiple tasks, M and D are represented as vector type, while Fnc, Lb, and Ub are represented as cell type. Notably, the function Fnc for each task takes the decision variable Dec as input and returns objective values Obj along with the constraint violation value CV. Moreover, for multi-objective optimization, an external function getOptimum() is required to define optimal solutions (i.e., reference points).
@@
With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP. The existing problems available in MToP are enumerated in Table VII. In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
TABLE VII: Multitask benchmark and real-world optimization problems included in MToP (current version).
Problem Case Task Dimension Objective Special Description
CEC17-MTSO [106] 9 2 50 Single - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MaTSO [107] 6 2-50 50 Single - Single-objective many-task optimization problems
WCCI20-MTSO [108] 10 2 50 Single - Complex single-objective multitask optimization problems
WCCI20-MaTSO [108] 10 2-50 50 Single - Complex single-objective many-task optimization problems
CMT [45] 9 2 Any Single Constrained Large/partial/no feasible regions interaction and different inter-task similarity
C2TOP&C4TOP [33] 51 2/4 50 Single Competitive Different optimal objective values among competitive tasks
C-CPLX [44] 20 2 50 Single Competitive Complex competitive multitask optimization problems
CEC17-MTMO [109] 9 2 50 Multi - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MTMO [107] 10 2 50 Multi - Complex multi-objective multitask optimization problems
CEC19-MaTMO [107] 10 2-50 50 Multi - Multi-objective many-task optimization problems
WCCI20-MaTMO [108] 10 2-50 50 Multi - Complex multi-objective many-task optimization problems
CEC21-MTMO [110] 10 2 50 Multi - More complex multi-objective multi optimization problems
PKACP [30] 1 Any Any Single - Planar kinematic arm control problem as many-task optimization problem
SCP [33] 1 Any 60-120 Single Competitive Sensor coverage problem as competitive multitask optimization problem
PEPVM [61] 1 3 7 Single - Parameter extraction of photovoltaic models as multitask optimization problem
OPF [59] 5 2 24/33 Single Constrained Optimal power flow as constrained multitask optimization problem
SOPM [69] 2 3 30 Multi Constrained Synchronous optimal pulse-width modulation as constrained multiobjective multitask problem
MO-OPF [69] 4 2 24/33 Multi Constrained Multiobjective optimal power flow as constrained multiobjective multitask problem
MO-SCP [61] 1 Any 60-120 Multi Competitive Multiobjective sensor coverage problem as competitive multiobjective multitask problem
@@ II-C3 Performance Metric
Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions. An illustrative example of metric implementation is presented in Listing LABEL:list:metric. The input parameter of the function is MTOData, which is generated during experimental execution. The function returns result, comprising properties such as Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData. The Metric property can take values of either Min’ or Max’, indicating whether a smaller or larger metric value is preferable. RowName, ColumnName, and TableData are utilized to present metric results in the GUI table. On the other hand, ConvergeData and ParetoData are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively. Note that specific metric calculation functions are not elaborated here for the sake of simplicity. The metrics currently integrated into MToP are enumerated in Table VIII.
TABLE VIII: Metrics included in MToP (current version).
@@
Obj Single All Objective value for each task
Obj (MTS) Single Multi Multitask score of Obj for all tasks
Obj (AV) Single Multi Average Obj for all tasks
Obj (CMT) Single Multi Competitive multitask Obj for all tasks
Obj (NBR) Single Multi Number of best Obj result for all tasks
CV Single All Constraint violation for each task
FR Single All Feasible rate for each task
HV Multi All Hypervolume for each task
HV (MTS) Multi Multi Multitask score of HV for all tasks
HV (CMT) Multi Multi Competitive multitask HV for all tasks
IGD Multi All Inverted generational distance for each task
IGD+ Multi All Improved IGD plus for each task
IGD+ (MTS) Multi Multi Multitask score of IGD+ for all tasks
IGD+ (CMT) Multi Multi Competitive multitask IGD+ for all tasks
Run Time All All Algorithm running time for all tasks
@@ III Performing Experiments with MToP
Figure 7: Graphical user interface of MToP.
To launch MToP, start by running the script file mto.m located in the root directory. This action will initialize the MToP GUI interface, which is illustrated in Fig. 7.
@@ III-A Testing Algorithms and Problems
As depicted in Fig. 7 (a), the test module interface is structured into left, center, and right sections. The left section facilitates the selection and configuration of algorithms and problems. Upon choosing the desired labels such as Task, Objective, and Special from the drop-down buttons, the options available for Algorithm and Problem are updated accordingly. Selecting specific algorithms and problems reveals their parameters in the edit box, which can be edited by clicking on them. Once parameters are adjusted, the new values are applied to the algorithm and problem objects through the setParameter() method. Algorithm parameters typically include evolutionary operator settings and other customizable parameters, while problem parameters usually consist of population size N, maximum function evaluations maxFE, and other specific parameters.
@@ III-B Executing Comparative Experiments
In the first column, users can configure the basic experiment settings. The No. of Runs textbox specifies the number of independent repeat runs for each algorithm on each problem. The No. of Results textbox determines the number of results saved during the evolutionary process of each algorithm, sampled uniformly according to the number of generations. The Parallel button toggles parallel processing, enabling multiple threads to run independent experiments simultaneously. The drop-down buttons Task, Objective, and Special filter algorithms and problems based on predefined labels, similar to the test module. Users can add selected algorithms and problems by clicking the Add button after choosing them from the respective drop-down boxes at the bottom.
In the second column, the names and parameters of the added algorithms and problems can be edited by clicking on them. Once all configurations are set, clicking the green Start button initiates the experiment execution. Users can also pause or stop the experiment using the Pause and Stop buttons if needed. Once the experiment is completed, the table’s contents become available in the results area on the right part of the experiment module. Users can read and save the original data of the experiment using the Load Data and Save Data buttons.
In the results area, the first function row contains buttons for operating on the experimental data. The first button, labeled Save Table, exports the table results to xlsx, csv, and tex files. The subsequent buttons labeled Converge control the convergence y-lim types and generate convergence plots based on the selected results from the table. The Pareto button draws the Pareto front and non-dominated solutions based on the selected results.
@@ III-C Executing via Alternative Command Line
In addition to GUI execution, MToP offers alternative options for conducting experiments via the command line. By inputting the algorithm list, problem list, and other parameters into the script file mto.m, experiments can be executed directly from the command line. Upon completion, the experiment data MTOData.mat is returned. Similarly, by inputting the data into the metrics function, users can obtain the calculated metric results.
@@ IV-A Adding New Algorithms
To incorporate a new algorithm into MToP, users should first generate a MATLAB class file within the Algorithms/ folder. This class file should inherit from the base class Algorithm.m. Subsequently, users can implement the algorithmic framework following the provided code pattern, as demonstrated in Listing LABEL:list:algo. Utilizing functions such as Initialization(), Generation(), and Selection() provided by MToP is recommended. However, users can also choose to reimplement these functions to suit their specific requirements. Additionally, users seeking to create customizable parameters that can be modified via the MToP GUI should implement the getParameter() and setParameter() APIs within the new algorithm class file.
@@ IV-B Adding New Problems
To integrate new problems into MToP, users should create a dedicated problem class file within the Problems/ folder. This class file should inherit from the base class Problem.m. It’s advisable to define the suggested maximum function evaluations (maxFE) within the constructor of the specific problem. Subsequently, users can refer to Listing LABEL:list:prob for guidance on configuring each optimization task within the setTasks() function. In cases where optimization tasks involve multi-objective optimization problems, users need to additionally implement the getOptimum() function to provide the optimal solution set or reference points. Lastly, for setting customizable parameters that can be modified within the GUI, users must also implement the getParameter() and setParameter() functions.
@@ IV-C Adding New Performance Metrics
To incorporate new performance metrics into MToP, users should create a new metrics function file within the Metrics/ folder. Follow the structure outlined in Listing LABEL:list:metric to receive the data structure object MTOData and return the calculated result object. For the metrics displayed in the GUI table, ensure that the RowName, ColumnName, and TableData fields are properly populated. Typically, the RowName for combined multitask metrics is the problem name, while metrics on individual tasks should return the name of each task. The ColumnName typically represents the algorithm name. Implement the calculation function for the specific performance metric in the TableData field. For plotting metric convergence, ensure the ConvergeData field is populated. In the case of multi-objective optimization, include non-dominated solutions in the ParetoData field.
@@ V Conclusion and Future Work
Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community. It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field. Specifically, MToP is considering the incorporation of multitask combinatorial optimization problems [5, 6] and more real-world applications [8, 111] in the future.

### v3

@@ II-C2 Problem
TABLE IV: Properties and methods of problem base class.
@@ D Number of decision variable dimensions for all tasks
getOptimum() Get optimal solutions for multi-objective optimization
@@ 11 Prob = Prob@Problem(varargin);
All problems within MToP inherit from the problem base class <Problem.m>. The properties and methods of <Problem.m> are detailed in Table IV. As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach [8]. It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
@@
A simple problem implementation example is illustrated in Listing LABEL:list:prob. The maximum number of function evaluations, denoted as <maxFE>, for the problem can be specified within the class constructor. The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task. Within this function, the number of tasks <T> is first defined, followed by a detailed setup for each task. Subsequently, parameters such as decision variable dimensions <D>, objective dimensions <M>, fitness function <Fnc>, lower bound <Lb>, and upper bound <Ub> are set individually for each task. For problems involving multiple tasks, <M> and <D> are represented as <vector> type, while <Fnc>, <Lb>, and <Ub> are represented as <cell> type. Notably, the function <Fnc> for each task takes the decision variable <Dec> as input and returns objective values <Obj> along with the constraint violation value <CV>. Moreover, for multi-objective optimization, an external function getOptimum() is required to define optimal solutions (i.e., reference points).
@@
With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP. The existing problems available in MToP are enumerated in Table LABEL:supp:tab:prob-mt of the supplementary files. In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
@@ II-C3 Performance Metric
Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions. An illustrative example of metric implementation is presented in Listing LABEL:list:metric. The input parameter of the function is <MTOData>, which is generated during experimental execution. The function returns <result>, comprising properties such as <Metric>, <RowName>, <ColumnName>, <TableData>, <ConvergeData>, and <ParetoData>. The <Metric> property can take values of either ’Min’ or ’Max’, indicating whether a smaller or larger metric value is preferable. <RowName>, <ColumnName>, and <TableData> are utilized to present metric results in the GUI table. On the other hand, <ConvergeData> and <ParetoData> are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively. Note that specific metric calculation functions are not elaborated here for the sake of simplicity. The metrics currently integrated into MToP are enumerated in Table V.
TABLE V: Metrics included in MToP. The table lists the metric name, optimization objective, task type, and a brief description. SO and MO denote single-objective and multi-objective optimization, respectively. ST and MT represent single-task and multi-task, respectively
@@
Obj SO ST/MT Objective value for each task
Obj (AV) SO MT Average Obj for all tasks
Obj (UV) SO MT Unified average Obj for all tasks
Obj (MTS) SO MT Multitask score of Obj for all tasks
Obj (CMT) SO MT Competitive multitask Obj for all tasks
Obj (NBR) SO MT Number of best Obj result for all tasks
CV SO ST/MT Constraint violation for each task
FR SO ST/MT Feasible rate for each task
HV MO ST/MT Hypervolume for each task
HV (MTS) MO MT Multitask score of HV for all tasks
HV (CMT) MO MT Competitive multitask HV for all tasks
IGD MO ST/MT Inverted generational distance for each task
IGD (AV) MO MT Average IGD for all tasks
IGD+ (MTS) MO MT Multitask score of IGD for all tasks
IGD (CMT) MO MT Competitive multitask IGD for all tasks
IGD+ MO ST/MT Improved IGD plus for each task
IGD+ (MTS) MO MT Multitask score of IGD+ for all tasks
IGD+ (CMT) MO MT Competitive multitask IGD+ for all tasks
Spread MO ST/MT Spread metric for each task
Spread (CMT) MO MT Competitive multitask spread for all tasks
Run Time SO/MO ST/MT Algorithm running time for all tasks
@@ III Performing Experiments with MToP
Figure 6: Graphical user interface of MToP. The test module (a) is used for testing algorithms and problems. The graphical display is shown in the center section of the test module. The experiment module (b) is used for executing comparative experiments. The results of the experiment are displayed in the right section of the experiment module.
To launch MToP, start by running the script file <mto.m> located in the root directory. This action will initialize the MToP GUI interface, which is illustrated in Fig. 6. The GUI interface of MToP requires MATLAB R2020b or later versions to run, while command-line execution can be done with any version. In addition, if parallel execution is required, MATLAB Parallel Computing Toolbox is needed.
@@ III-A Testing Algorithms and Problems
As depicted in Fig. 6 (a), the test module interface is structured into left, center, and right sections. The left section facilitates the selection and configuration of algorithms and problems. Upon choosing the desired labels such as <Task>, <Objective>, and <Special> from the drop-down buttons, the options available for <Algorithm> and <Problem> are updated accordingly. Selecting specific algorithms and problems reveals their parameters in the edit box, which can be edited by clicking on them. Once parameters are adjusted, the new values are applied to the algorithm and problem objects through the setParameter() method. Algorithm parameters typically include evolutionary operator settings and other customizable parameters, while problem parameters usually consist of population size <N>, maximum function evaluations <maxFE>, and other specific parameters.
@@ III-B Executing Comparative Experiments
In the first column, users can configure the basic experiment settings. The <No. of Runs> textbox specifies the number of independent repeat runs for each algorithm on each problem. The <No. of Results> textbox determines the number of results saved during the evolutionary process of each algorithm, sampled uniformly according to the number of generations. The <Parallel> button toggles parallel processing, allowing independent experiments to run simultaneously across multiple process. The number of concurrent experiments depends on the number of processor cores. If the number of parallel experiments exceeds the available cores, MATLAB automatically handles the scheduling. For multiple algorithms and problems, each algorithm processes a problem in parallel. After one algorithm finishes, the next starts, repeating until all algorithms have processed the problem, then moving to the next problem. Hence, the number of algorithms and problems is not inherently limited. The drop-down buttons <Task>, <Objective>, and <Special> filter algorithms and problems based on predefined labels, similar to the test module. Users can add selected algorithms and problems by clicking the <Add> button after choosing them from the respective drop-down boxes at the bottom.
In the second column, the names and parameters of the added algorithms and problems can be edited by clicking on them. Once all configurations are set, clicking the green <Start> button initiates the experiment execution. Users can also pause or stop the experiment using the <Pause> and <Stop> buttons if needed. Once the experiment is completed, the table’s contents become available in the results area on the right part of the experiment module. Users can read and save the original data of the experiment using the <Load Data> and <Save Data> buttons.
In the results area, the first function row contains buttons for operating on the experimental data. The first button, labeled <Save Table>, exports the table results to xlsx, csv, and tex files. The subsequent buttons labeled <Converge> control the convergence y-lim types and generate convergence plots based on the selected results from the table. The <Pareto> button draws the Pareto front and non-dominated solutions based on the selected results.
@@ III-C Executing via Alternative Command Line
In addition to GUI execution, MToP offers alternative options for conducting experiments via the command line. By inputting the algorithm list, problem list, and other parameters into the script file <mto.m>, experiments can be executed directly from the command line. For example, running the command mto({MFEA, EMEA}, {CMT1, CMT2}, 30) will execute the MFEA and EMEA algorithms on the CMT1 and CMT2 problems independently 30 times. Upon completion, the experiment data <MTOData.mat> is returned. Similarly, by inputting the data into the metrics function, users can obtain the calculated metric results.
@@ IV-A Adding New Algorithms
To incorporate a new algorithm into MToP, users should first generate a MATLAB class file within the <Algorithms/> folder. This class file should inherit from the base class <Algorithm.m>. Subsequently, users can implement the algorithmic framework following the provided code pattern, as demonstrated in Listing LABEL:list:algo. Utilizing functions such as Initialization(), Generation(), and Selection() provided by MToP is recommended. However, users can also choose to reimplement these functions to suit their specific requirements. Additionally, users seeking to create customizable parameters that can be modified via the MToP GUI should implement the getParameter() and setParameter() APIs within the new algorithm class file.
@@ IV-B Adding New Problems
To integrate new problems into MToP, users should create a dedicated problem class file within the <Problems/> folder. This class file should inherit from the base class <Problem.m>. It’s advisable to define the suggested maximum function evaluations <maxFE> within the constructor of the specific problem. Subsequently, users can refer to Listing LABEL:list:prob for guidance on configuring each optimization task within the setTasks() function. In cases where optimization tasks involve multi-objective optimization problems, users need to additionally implement the getOptimum() function to provide the optimal solution set or reference points. Lastly, for setting customizable parameters that can be modified within the GUI, users must also implement the getParameter() and setParameter() functions.
@@ IV-C Adding New Performance Metrics
To incorporate new performance metrics into MToP, users should create a new metrics function file within the <Metrics/> folder. Follow the structure outlined in Listing LABEL:list:metric to receive the data structure object <MTOData> and return the calculated <result> object. For the metrics displayed in the GUI table, ensure that the <RowName>, <ColumnName>, and <TableData> fields are properly populated. Typically, the <RowName> for combined multitask metrics is the problem name, while metrics on individual tasks should return the name of each task. The <ColumnName> typically represents the algorithm name. Implement the calculation function for the specific performance metric in the <TableData> field. For plotting metric convergence, ensure the <ConvergeData> field is populated. In the case of multi-objective optimization, include non-dominated solutions in the <ParetoData> field.
IV-D Extending Optimization Types
There are also some areas, such as sequential transfer optimization and multiform optimization, in the evolutionary transfer optimization field similar to evolutionary multitasking [18]. Although MToP is currently focused on evolutionary multitasking, it is possible to extend the platform to include these new optimization types. Users can refer to the existing code patterns in MToP to develop new optimization types. Specifically, for the sequential transfer optimization, it is possible to set the target task as the first task and the remaining tasks as source tasks for algorithms and problems. The metric calculation can be adjusted to evaluate the performance of the target task. For multiform optimization, it can dynamically construct multiform tasks within the implemented algorithms depending on the problem to be solved.
@@ V Conclusion and Future Work
Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community. It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field. Specifically, MToP is considering the incorporation of many-objective optimization codes [46, 4], multitask combinatorial optimization codes [5, 6] and more real-world applications [9, 47] in the future.